Solve and check the linear equation 2x-2=18


(4x -2)+7= -3(x+3)

Answer by checkley77(12844)

2x-2=18
2X=18+2
2X=20
X=20/2
X=10 ANS.
-------------------------
(4x -2)+7= -3(x+3)
4X-2+7=-3X-9
4X+3X=-9-7+2
7X=-14
X=-14/7
X=-2 ANS.